Find the measures of one of the exterior angles at the base of an isosceles triangle if one of the angles of the triangle is 80 degrees??
Answer by richwmiller(17219)

an isosceles triangle has two angles equal
we don't know if there are two 80 degree angles or just one
so we have two options
first
80+80=160
180-160=20
angles 80,80,20
180-80=100 exterior angle
second
180-80=100
100/2=50
angles 50,50,80
180-50=130 exterior